You are given the following information regarding prices for a sample of stocks.

PRICE
Stock Number of Shares T T + 1
A   1,000,000 $68 $88
B 10,000,000   30   42
C 25,000,000   25   37
  1. Construct a price-weighted index for these three stocks, and compute the percentage change in the index for the period from T to T + 1. Do not round intermediate calculations. Round your answer to two decimal places.

      %

  2. Construct a value-weighted index for these three stocks, and compute the percentage change in the index for the period from T to T + 1. Do not round intermediate calculations. Round your answer to two decimal places.

      %
Answer #1

a).

Stock Period T Period T+1
A 68 88
B 30 42
C 25 37
Sum 123 167
Divisor 3 3
Average = Sum/Divisor 41 55.67

% change = [Period (T+1)'s Average - Period T's Average] / Period T's Average

= [55.67 - 41] / 41 = 14.67 / 41 = 0.3577, or 35.77%

b). For Period 'T':

Stock Share Price Number of Shares Market Value
A $68 1,000,000 $68,000,000
B $30 10,000,000 $300,000,000
C $25 25,000,000 $625,000,000
Total $993,000,000

For Period 'T+1':

Stock Share Price Number of Shares Market Value
A $88 1,000,000 $88,000,000
B $42 10,000,000 $420,000,000
C $37 25,000,000 $925,000,000
Total $1,433,000,000

% change = [Period (T+1)'s Average - Period T's Average] / Period T's Average

= [$1,433,000,000 - $993,000,000] / $993,000,000

= $440,000,000 / $993,000,000 = 0.4431, or 44.31%